LHC is now in technical stop (week 51) and then Christmas break until 7th January 2013.
Proton-lead run scheduled to start 14th January.
Goal of this week: 25 ns scrubbing, 25 ns ramp and squeeze for MD and physics run. 3 days contingency at end of the week.
Link to report from morning meeting
|Mon 17 Dec
- longitudinal damper studies and instability monitor tests at 450 GeV
- quick attempt for physics with 800b 25 ns - dumped by ALICE going into collisions - luminosity too high
- 06:00 Ramp down and prepare for technical stop
Summary of week's e-cloud studies etc (Giovanni Rumolo)
|Sun 16 Dec
- 00:00 new injection: 12+4*96 bunches, but problem with timing: 3 batches out of time by 10ns ... due to problems in optic fiber length compensation, -> beam dump
- 06:40 timing problem solved
- 09:00 injection probems: RF clock hangover issues
- Fill for 396b 25 ns fill - no problem (although e-cloud induced blow-up observed)
- 21:45 to adjust for short collimator EOF study
|Sat 15 Dec
- Patrol lost in a few sectors in point 6
- Ready for injection at 03:00 Inject trains of 288 bunches.
- Dumped by injection losses in point 2 at 05:10 when we had 1884 bunch/beam. Heat load not yet stabilized.
- Fill the machine with trains of 72 bunches (2100 bunches with a stop at ~1000 bunches) for benchmark purposes. Let the heat load to stabilize.
- During filling (06:30) problem with TDI in point 2. OK after piquet intervention.
- 15:00 injection 12+2*48 bunches in h9 mode, ramp, squeeze adjust &stable beams
- 20:00 injection 12+4*48 bunches in h9 brought into stable beams
|Fri 14 Dec
- 804 bunches for e-cloud conditioning at flat top
- Waiting for a repair of the RBIH400107 power converter in TT40.
- Verification of injection timing of h=9 beam with TEDs in
- 11:00 Repair of TT40 power supply Fill for loss maps – 12b
- 12:30 loss maps end of squeeze
- 13:10 loss maps in collision
- Restarted for 25 ns MD at 16:30 after Loss maps and h=9 injection test
- Injection of 804 bunches with reduced bunch intensity (0.9x1011 p/b) for heat load comparison at different bunch intensity for benchmarking purposes
- First ramp lost due to SW interlock (several BPM not returning data)
- Difficulties in filling again (octupoles not set correctly, unexplained BPM trigger at injection)
- 21:50 Finally at flat-top after an UFO (67% of dump threshold) during the ramp
- 23:04 beams dumped due to loss of communication with a few access elements in point 6
25 ns MD summary
|Thu 13 Dec
- Wrestling match with RF
- Final get up ramp, through squeeze and collide 25 ns beams
- 06:15 Lumi peak at ~83e30 cm-2s-1 With 9e12 p per ring, 72 colliding bunches/ring and 12 non colliding per ring, that gives an emittance at the start of collisions of ~3.8 um.
- Start long range MD
- 08:30 once more: loss maps at flat top to varify collimator settings before intensty fills
- 11:50 injection & ramp: 12+2*72 bunches per beam
- 15:00 injection & ramp: 12+5*72 bunches per beam
- 20:00 injection &ramp: 12+11*72 bunches per beam
- kept in flat top for e-cloud conditioning
|Wed 12 Dec
- 07:00 test ramp from injection to squeeze and measure emittances
- 10:00 TDI alignment checks on beam
- ~16:00 TDI checks finished (some delays injector side)
- 18:30 MKI-B2 didn't pulse while attempting to inject 48b (12b circulating)
- 19:30 B2 48b went again onto the TDI and the circulating 12b were dumped. MKI8 pulsed… To be following up today (CO/RF/ABT/BIS…)
- 19:30 Switch to standard 72*25 ns
- 20:03 RF modules M1B1 and M2B2 tripped while increasing power supply voltage from 50kV->58kV
|Tues 11 Dec
- 02:50 Flat-top with 3 nominals
- 03:30 Looking for collisions. Incorporated the lumi scans at the knob level and the orbit H and V at the K level.
- 05:00 Beams dumped Back around again with 3 nominals for collimator set-up
- 08:00 Beams colliding – collimator team into action (Stefano and Gianluca)… ongoing
- 11:30 short loss of cryo due to error in T-reading
- 12:00 access for QPS FGC and TDI problem
- 18:00 injection of 4 bunches for collimator alignement & loss maps to validate
|Mon 10 Dec
- 25 ns fills for scrubbing at 450 GeV
- 08:22 Access for ATLAS, QPS and TDI point 8.
- 13:00 Test ramp, squeeze and collide for 25 ns squeeze to 1 m with relaxed collimator settings
- 15:50 Test ramp with nominals - beams lost on flat-top - cryogenics problem
- 23:00 Cryo-start re-established (with some reservations) - warm front still propagating through the arc - start cycling